OpenStudy (anonymous):

For which equation would x = 1 not be a solution? x/1 = 1 x/2 = 2/4 x/3 = 3 x/4 = 1/4

OpenStudy (anonymous):

i think the second one

OpenStudy (anonymous):

just plug in what x is :)

OpenStudy (anonymous):

hi some_someone

OpenStudy (anonymous):

x = 1 x/1 = 1 1/1 = 1 TRUE

OpenStudy (anonymous):

thanks i got more so be ready

OpenStudy (anonymous):

x = 1 x/2 = 2/4 1/2 = 2/4 1/2 = 1/2 TRUE

OpenStudy (anonymous):

can u do it?

OpenStudy (anonymous):

x = 1 x/3 = 3 1/3 = 3 FALSE

OpenStudy (anonymous):

wait u can do anything

OpenStudy (anonymous):

x = 1 x/4 = 1/4 1/4 = 1/4 TRUE

OpenStudy (anonymous):

so second one or 4th one

OpenStudy (anonymous):

which one is FALSE ?

OpenStudy (anonymous):

the third one

OpenStudy (anonymous):

precisely :)

OpenStudy (anonymous):

Therefore, x = 1 could not be a solution for x/3 = 3